Circular No. 8777

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 W3 (CHRISTENSEN) E. J. Christensen reports his discovery of a comet on Catalina Sky

Survey images obtained in good seeing with the 0.68-m Schmidt telescope (discovery observation tabulated below); four 60-s exposures showed a 15′′ coma extended very slightly toward p.a. 70^{o}–100^{o}. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, several astrometric CCD observers have commented on the object’s cometary appearance. Stacked 15-s images obtained by L. Buzzi and F. Luppi (Varese, Italy, 0.60-m f/4.6 reflector) on Nov. 19.1 UT show a 10′′ coma elongated in p.a. 80^{o} with a 25′′ tail in the same direction. P. Birtwhistle (Great Shefford, Berkshire, England, 0.40-m f/6 Schmidt-Cassegrain reflector) writes that his images from Nov. 19.1 show the object to be distinctly diffuse and only moderately condensed, with the center of light slightly offset toward the west in a 14′′ coma (no tail). J. McGaha (Tucson, AZ, 0.62-m f/5.1 reflector) reports that his 120-s exposures on Nov. 19.4 show a 9′′ coma elongated in p.a. 35^{o}, with no apparent nuclear condensation or tail. J. G. Ries (McDonald Observatory, 0.76-m reflector) notes that his R-band images on Nov. 19.4 show an asymmetric diffuse coma that is more prominent on the side opposite to the comet’s direction of motion.

2006 UT  α_{2000}  δ_{2000}  Mag.
Nov. 18.40414  7^{h}00^{m}04.^{s}24  +46^{o}16′11′′.3  18.1

The available astrometry, the following very preliminary orbital elements, and an ephemeris appear on MPEC 2006-W42.

T = 2008 Dec. 25.327 TT  ω = 154.^{o}652
Ω = 109.832 }2000.0
q = 1.46888 AU  i = 121.736

COMET C/2006 U7 (GIBBS) Additional astrometry of this comet (cf. IAUC 8769) reveals an ellip-

tical orbit (the following elements taken from MPEC 2006-W18):

T = 2007 Feb. 10.974 TT  ω = 6.^{o}897
e = 0.63310  Ω = 57.906 }2000.0
q = 4.40745 AU  i = 6.992
a = 12.01272 AU n^{o} = 0.023672 P = 41.6 years
2006 November 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green